Please help. I've been working on trying to solve by setting up a polynomial expression, but having trouble with this one.
If the sides of a square are decreased by 2 cm, the area is decreased by 36 cm^2. What were the dimensions of the original square?
Thanks in advance.

Answer by edjones(8007)

let x=length of original side.
A=s^2
new square area=old square area-36
(x-2)^2=x^2-36
x^2-4x+4=x^2-36
subtract x^2 and 4 from both sides: -4x=-40
divide -4 into each side: x=10 Ans.
Check:
10^2=100
8^2=64
100-64=36
Ed